OpenStudy (anonymous):

which equation represents y=-x^2-2x+3 in vertex form?

OpenStudy (anonymous):

help needed?

OpenStudy (apoorvk):

Well, this is a maths question, request you kindly post questions in the right forum next time onwards :)

OpenStudy (apoorvk):

you need complete the square of the quadratic in 'x'. x^-2x+3 = x^2 -2x + 1 +2 = (x-1)^2 + 2 --> = y or, y-2 = (x-1)^2 Now here, the quadratic of the form (y+k)=(x+h)^2 has its vertex at (-h,-k).